OpenStudy (anonymous):

Simplify the expression (4x − 3)(x + 5).

OpenStudy (anonymous):

4*x^2-3*x+20*x-15 =4*x^2+17*x-15

OpenStudy (anonymous):

use FOIL :)

OpenStudy (anonymous):

It's simple multiplication, first, multiply 4x-3 with x, then with 5 and add.

OpenStudy (anonymous):

(4x)(x)+(4x)(5)-(3)(x)-(3)(5)=? simplify :) combine like terms :D

OpenStudy (anonymous):

(4x-3)(x+5) \[4x ^{2}+20x-3x-15\] \[4x ^{2}+17x-15\]

OpenStudy (anonymous):

its like (a+b)(a+c) = a(a+c)+b(a+c) \[a ^{2}+ac+ba+bc\]

OpenStudy (anonymous):

I THINK ITS BETTER IF YOU ALSO LET HIM/HER DO HIS PART :/ TSK ITS THE MATTER OF TEACHING HIM, NOT GIVING THE ANSWER RIGHT AWAY